# GenMol: A Drug Discovery Generalist with Discrete Diffusion

### v1

@@ Abstract
Abstract: Drug discovery is a complex process that involves multiple scenarios and stages, such as fragment-constrained molecule generation, hit generation and lead optimization. However, existing molecular generative models can only tackle one or two of these scenarios and lack the flexibility to address various aspects of the drug discovery pipeline. In this paper, we present Generalist Molecular generative model (GenMol), a versatile framework that addresses these limitations by applying discrete diffusion to the Sequential Attachment-based Fragment Embedding (SAFE) molecular representation. GenMol generates SAFE sequences through non-autoregressive bidirectional parallel decoding, thereby allowing utilization of a molecular context that does not rely on the specific token ordering and enhanced computational efficiency. Moreover, under the discrete diffusion framework, we introduce fragment remasking, a strategy that optimizes molecules by replacing fragments with masked tokens and regenerating them, enabling effective exploration of chemical space. GenMol significantly outperforms the previous GPT-based model trained on SAFE representations in de novo generation and fragment-constrained generation, and achieves state-of-the-art performance in goal-directed hit generation and lead optimization. These experimental results demonstrate that GenMol can tackle a wide range of drug discovery tasks, providing a unified and versatile approach for molecular design.
@@ 1 Introduction
Discovering molecules with the desired chemical profile is the core objective of drug discovery (Hughes et al., 2011). To achieve the ultimate goal of overcoming disease, a variety of drug discovery approaches have been established. For example, fragment-constrained molecule generation is a popular strategy for designing new drug candidates under the constraint of preserving a certain molecular substructure already known to exhibit a particular bioactivity (Murray & Rees, 2009). Furthermore, real-world drug discovery pipelines are not a single stage but consist of several key stages, such as hit generation and lead optimization (Hughes et al., 2011). A drug discovery process that leads to the finding of successful drug candidates that can enter clinical trials should consider all of these different scenarios.
Figure 1: Results on various drug discovery tasks. The values are quality, average quality, sum AUC top-10, and success rate for de novo generation, fragment-constrained generation, hit generation, and lead optimization, respectively. The “best baseline” refers to multiple best-performing task-specific models among prior works.
@@
Generative models have emerged as a promising methodology to greatly accelerate labor-intensive drug discovery pipelines (Olivecrona et al., 2017; Jin et al., 2018; Yang et al., 2021; Lee et al., 2023), but previous molecular generative models have a common limitation: they focus on only one or two of the drug discovery scenarios. They either cannot be applied to multiple tasks or require expensive modifications including retraining of a specific architecture for each task (Yang et al., 2020; Guo et al., 2023). Recently, SAFE-GPT (Noutahi et al., 2024) has been proposed to address this problem by using Sequential Attachment-based Fragment Embedding (SAFE) that represents a molecule as an unordered sequence of Simplified Molecular Input Line Entry System (SMILES) (Weininger, 1988) fragment blocks and viewing fragment-constrained generation as a sequence completion task. However, since all sequence-based molecular representations including SMILES and SAFE have an ordering of tokens based on heuristic rules such as depth-first search (DFS), autoregressive models that operate in a left-to-right order like GPT are unnatural for processing and generating molecular sequences. In addition, the autoregressive decoding scheme limits the computational efficiency of the model. Moreover, SAFE-GPT relies on finetuning under expensive reinforcement learning (RL) objectives to be applied to goal-directed molecule generation (Noutahi et al., 2024).
To this end, we propose Generalist Molecular generative model (GenMol), a versatile molecular generation framework that is endowed with the ability to handle diverse scenarios that can be encountered in the multifaceted drug discovery pipeline (Figure 2). Specifically, GenMol adopts a masked discrete diffusion framework (Austin et al., 2021; Sahoo et al., 2024) with the BERT architecture (Devlin et al., 2019) to generate SAFE molecular sequences, thereby enjoying several advantages. Discrete diffusion allows GenMol to exploit a molecular context that does not rely on the specific ordering of tokens and fragments by bidirectional attention. Additionally, the non-autoregressive parallel decoding improves GenMol’s computational efficiency without degrading the generation quality (Figure 2(b)). Furthermore, discrete diffusion enables GenMol to explore chemical space with a simple yet effective remasking strategy. Concretely, we propose fragment remasking (Figure 2(c6) and Figure 3), a strategy to optimize molecules by replacing certain fragments in a given molecule with masked tokens, from which discrete diffusion generates new fragments. By utilizing fragments as the explorative unit instead of individual tokens, GenMol can effectively and efficiently explore the vast chemical space to find chemical optima.
@@
  • We introduce GenMol, a framework for unified and versatile molecule generation through applying discrete diffusion to the SAFE representation.
@@ 2 Related Work
There has been steady progress in applying discrete diffusion for discrete data generation, especially in NLP tasks (Hoogeboom et al., 2021; Austin et al., 2021; He et al., 2023; Zheng et al., 2024; Lou et al., 2024; Sahoo et al., 2024). This is mainly due to their non-autoregressive generation property, which leads to a potential for better modeling long-range dependencies and acclerating sampling speed, and their flexible design choices in training, sampling, and controllable generation (Sahoo et al., 2024). Notably, D3PM (Austin et al., 2021) introduced a general framework with a Markov forward process represented by transition matrices, and a transition matrix with an absorbing state corresponds to the masked language modeling (MLM) such as BERT (Devlin et al., 2019). Campbell et al. (2022) proposed a continuous-time framework for discrete diffusion models based on the continuous-time Markov chain (CTMC) theory. SEDD (Lou et al., 2024) introduced a denoising score entropy loss that extends score matching to discrete diffusion models. Sahoo et al. (2024) and Shi et al. (2024) proposed simple masked discrete diffusion frameworks, with the training objective being a weighted average of MLM losses across different diffusion time steps.
@@
Fragment-based molecular generative models refer to a class of methods that reassemble existing molecular substructures (i.e., fragments) to generate new molecules. A line of works (Jin et al., 2020; Maziarz et al., 2021; Kong et al., 2022; Geng et al., 2023) used graph-based VAEs to generate novel molecules conditioned on discovered substructures. Besides, Xie et al. (2020) proposed to progressively add or delete fragments of molecular graphs using Markov chain Monte Carlo (MCMC) sampling. Yang et al. (2021) proposed a reinforcement learning (RL) framework with prioritized experience replay to assemble fragments.
Graph-based genetic algorithms (GAs) (Jensen, 2019; Tripp & Hernández-Lobato, 2023) remain as a strong approach, which decompose parent molecules into fragments that are combined to generate an offspring molecule. Since their fragment-based strategies are limited by the fact that their generation is from random combinations of existing fragments with a local mutation of a small probability, they suffer from limited exploration in the chemical space. More recently, $f$-RAG (Lee et al., 2024a) introduced a fragment-level retrieval framework that augments the pre-trained molecular language model SAFE-GPT (Noutahi et al., 2024), where retrieving fragments from dynamically updated fragment vocabulary largely improves the exploration-exploitation trade-off. However, $f$-RAG (Lee et al., 2024a) still needs to train an information fusion module before adapting to various goal-oriented generation tasks.
@@ 3.1 Masked Diffusion
Among various discrete diffusion frameworks, we follow MDLM (Sahoo et al., 2024) to define our masked diffusion model, due to its simplicity and effectiveness.
Formally, we define $\boldsymbol{x}$ as a sequence of $L$ discrete tokens, each of which, denoted as $\boldsymbol{x}^{l}$, is a one-hot vector with $K$ categories (i.e., $\boldsymbol{x}^{l}_{i}\in\{0,1\}^{K}$ and $\sum_{i=1}^{K}\boldsymbol{x}^{l}_{i}=1$). Without loss of generality, we assume the $K$-th category represents the masking token, whose one-hot vector is denoted by $\bm{m}$ (i.e., $\bm{m}_{K}=1$). We also define $\text{Cat}(\cdot;\boldsymbol{\pi})$ as a categorical distribution with a probability $\boldsymbol{\pi}\in\Delta^{K}$, where $\Delta^{K}$ represents the simplex over $K$ categories.
The forward masking process in the masked diffusion independently interpolates between each token in clean data sequence $\boldsymbol{x}^{l}$ and the masking token $\bm{m}$, which is defined as
$q(\boldsymbol{z}^{l}_{t}|\boldsymbol{x}^{l})=\text{Cat}(\boldsymbol{z}^{l}_{t};\alpha_{t}\boldsymbol{x}^{l}+(1-\alpha_{t})\bm{m}),$ (1)
@@
$p_{\theta}(\boldsymbol{z}^{l}_{s}|\boldsymbol{z}^{l}_{t})=\begin{cases}\text{Cat}(\boldsymbol{z}^{l}_{s};\boldsymbol{z}^{l}_{t}),\quad\boldsymbol{z}^{l}_{t}\neq\bm{m},\\ \text{Cat}(\boldsymbol{z}^{l}_{s};\frac{(1-\alpha_{s})\bm{m}+(\alpha_{s}-\alpha_{t})\boldsymbol{x}^{l}_{\theta}(\boldsymbol{z}_{t},t)}{1-\alpha_{t}}),\,\boldsymbol{z}^{l}_{t}=\bm{m},\end{cases}$ (2)
where $\text{Cat}(\boldsymbol{x};\boldsymbol{p})$ is a categorical distribution over the vector $\boldsymbol{x}$ with probabilities given by the vector $\boldsymbol{p}$ and $\boldsymbol{x}_{\theta}(\boldsymbol{z}_{t},t)$ is a denoising network that takes the full $L$ tokens in the noisy data sequence $\boldsymbol{z}_{t}$ as input and predicts $L$ probability vectors for the clean data sequence. This parameterization designs the reverse process such that it does not change unmasked tokens. To train the denoising network $\boldsymbol{x}_{\theta}(\boldsymbol{z}_{t},t)$, the training objective, which implicitly approximates the negative ELBO (Sohl-Dickstein et al., 2015), is given by
@@ 3.2 SAFE Molecular Representation
Sequential Attachment-based Fragment Embedding (SAFE) (Noutahi et al., 2024) has recently been proposed to alleviate this problem. SAFE represents molecules as an unordered sequence of fragment blocks, thereby casting molecular design tasks into sequence completion tasks. SAFE is a non-canonical SMILES that restricts the arrangement of SMILES tokens corresponding to the same molecular fragment to be consecutive. Molecules are decomposed into fragments by the BRICS algorithm (Degen et al., 2008) and the fragments are concatenated using a dot token (“.”) while preserving their attachment points. SAFE is permutation-invariant on fragments, i.e., the order of fragments within a SAFE string does not change the molecular identity.
@@ 4 Method
Figure 3: (a) Goal-directed hit generation and lead optimization process with GenMol. An initial fragment vocabulary is constructed by decomposing an existing molecule dataset (hit generation) or a seed molecule (lead optimization), two fragments are randomly sampled from the vocabulary and connected to provide a starting molecule to GenMol. GenMol performs fragment remasking, a strategy to modify molecules by randomly selecting a fragment in a given molecule and replacing it with a mask chunk, and then generating a new fragment. The fragment vocabulary is updated with the generated molecules for the next iteration. (b) Illustration of the molecular optimization trajectory with fragment remasking. With discrete diffusion-based fragment remasking, GenMol can explore beyond the initial fragment vocabulary to find chemical optima.
Now we introduce GenMol, a universal molecule generation framework that can solve various drug discovery tasks. We first introduce the construction of the discrete diffusion framework on the SAFE representation in Section 4.1. Next, we describe fragment remasking, an effective exploration strategy of GenMol for goal-oriented generation tasks, in Section 4.2.
@@ 4.1 Masked Diffusion for Molecule Generation
We apply discrete diffusion to the SAFE molecular representation to establish a flexible and efficient molecule generation framework. Specifically, GenMol adopts the architecture of BERT (Devlin et al., 2019) as the denoising network $\boldsymbol{x}_{\theta}$ and the training scheme of MDLM described in Section 3.1. Utilizing a discrete diffusion framework instead of an autoregressive model is more in line with the SAFE molecular representation and has several advantages. First, due to the bidirectional attention mechanism in BERT, GenMol can leverage parallel decoding where all tokens are decoded simultaneously under discrete diffusion (Figure 5). As SAFE sequences are fragment order-insensitive, this allows GenMol to predict masked tokens with a better inductive bias by looking backward and forward in the sequence. The non-autoregressive parallel decoding scheme also improves GenMol’s efficiency without degrading its generation quality. Furthermore, the discrete diffusion framework enables GenMol to explore the neighborhood of a given molecule in chemical space with a simple yet effective remasking strategy, which will be described in detail in Section 4.2.
During inference, inspired by Chang et al. (2022) and Tang et al. (2022), GenMol employs confidence sampling, a strategy that decides which tokens to unmask at each sampling step based on the confidence scores of the tokens. At each masked index $l$, GenMol samples a token $\boldsymbol{z}^{l}_{s}$ based on the reverse unmasking process $p_{\theta}(\boldsymbol{z}^{l}_{s}|\boldsymbol{z}_{t})$ calculated by $\boldsymbol{x}^{l}_{\theta}(\boldsymbol{z}_{t},t)$. Here, we introduce the softmax temperature $\tau$ to the logits of $\boldsymbol{x}^{l}_{\theta}(\boldsymbol{z}_{t},t)$ as follows:
$p^{l}_{\theta,i}=\frac{\exp\left(\log\boldsymbol{x}^{l}_{\theta,i}(\boldsymbol{z}^{l}_{t},t)/\tau\right)}{\sum^{K}_{i=1}\exp\left(\log\boldsymbol{x}^{l}_{\theta,i}(\boldsymbol{z}^{l}_{t},t)/\tau\right)}\text{ for}i=1,\cdots,K.$ (4)
After sampling from the probability specified by Eq. (4), the corresponding prediction score $p^{l}_{\theta,i^{*}}$ is used as the confidence score of $\boldsymbol{z}^{l}_{s}$, where $i^{*}$ is the index of the sampled token. We further introduce a gumbel noise decreasing over the sampling process, yielding the final confidence score of $\boldsymbol{z}^{l}_{s}$ as follows:
$\log p^{l}_{\theta,i^{*}}+r\cdot t\cdot\epsilon,\quad\epsilon\sim\text{Gumbel}(0,1),$ (5)
where $r$ is the randomness and $t$ is the current time step. Trade-offs between molecular quality and diversity often arise in drug discovery problems, and GenMol is able to balance these two essential considerations through the softmax temperature $\tau$ and the randomness $r$, as we will show in Section 5.1.
The number of tokens $N$ to unmask at each time step is defined by the mask schedule, and GenMol unmask the top-$N$ currently masked tokens based on their confidence scores (Eq. (5)). In other words, this is equivalent to predicting all tokens simultaneously, but only confirming the most confident predictions. The other tokens are masked again and predicted in the next step. With confidence sampling, GenMol can exploit the dependencies between tokens in a given molecular sequence for better sampling quality, rather than randomly and independently selecting tokens to unmask, as in the standard diffusion sampling (Austin et al., 2021; Sahoo et al., 2024).
4.2 Exploration with Fragment Remasking

### v2

@@ Abstract
Abstract: Drug discovery is a complex process that involves multiple stages and tasks. However, existing molecular generative models can only tackle some of these tasks. We present Generalist Molecular generative model (GenMol), a versatile framework that uses only a single discrete diffusion model to handle diverse drug discovery scenarios. GenMol generates Sequential Attachment-based Fragment Embedding (SAFE) sequences through non-autoregressive bidirectional parallel decoding, thereby allowing the utilization of a molecular context that does not rely on the specific token ordering while having better sampling efficiency. GenMol uses fragments as basic building blocks for molecules and introduces fragment remasking, a strategy that optimizes molecules by regenerating masked fragments, enabling effective exploration of chemical space. We further propose molecular context guidance (MCG), a guidance method tailored for masked discrete diffusion of GenMol. GenMol significantly outperforms the previous GPT-based model in de novo generation and fragment-constrained generation, and achieves state-of-the-art performance in goal-directed hit generation and lead optimization. These results demonstrate that GenMol can tackle a wide range of drug discovery tasks, providing a unified and versatile approach for molecular design.
@@ 1 Introduction
Discovering molecules with the desired chemical profile is the core objective of drug discovery (Hughes et al., 2011). To achieve the ultimate goal of overcoming disease, a variety of drug discovery approaches have been established. For example, fragment-constrained molecule generation is a popular strategy for designing new drug candidates under the constraint of preserving a certain molecular substructure already known to exhibit a particular bioactivity (Murray & Rees, 2009). Furthermore, real-world drug discovery pipelines are not a single stage but consist of several key stages, such as hit generation and lead optimization (Hughes et al., 2011). A drug discovery process that leads to the finding of drug candidates that can enter clinical trials should consider all of these different scenarios.
Figure 1: Results on drug discovery tasks. The values are quality, average quality, sum AUC top-10, and success rate for de novo generation, fragment-constrained generation, hit generation, and lead optimization, respectively. The “best baseline” refers to multiple best-performing task-specific models among prior works.
@@
Generative models have emerged as a promising methodology to accelerate labor-intensive drug discovery pipelines (Olivecrona et al., 2017; Jin et al., 2018; Yang et al., 2021; Lee et al., 2023), but previous molecular generative models have a common limitation: they focus on only one or two of the drug discovery scenarios. They either cannot be applied to multiple tasks or require expensive modifications including retraining of a specific architecture for each task (Yang et al., 2020; Guo et al., 2023). Recently, SAFE-GPT (Noutahi et al., 2024) has been proposed to address this problem by formulating several molecular tasks as a fragment-constrained generation task, solved by sequence completion. SAFE-GPT uses Sequential Attachment-based Fragment Embedding (SAFE) of molecules, which represents a molecule as an unordered sequence of Simplified Molecular Input Line Entry System (SMILES) (Weininger, 1988) fragment blocks. However, since all sequence-based molecular representations including SMILES and SAFE assume an ordering of tokens based on heuristic rules such as depth-first search (DFS), autoregressive models that operate in a left-to-right order like GPT are unnatural for processing and generating molecular sequences. In addition, the autoregressive decoding scheme limits the computational efficiency of the model and makes it challenging to introduce guidance during generation. Moreover, SAFE-GPT relies on finetuning under expensive reinforcement learning (RL) objectives to be applied to goal-directed molecule generation.
To tackle these limitations, we propose Generalist Molecular generative model (GenMol), a versatile molecular generation framework that is endowed with the ability to handle diverse scenarios that can be encountered in the multifaceted drug discovery pipeline (Figure 2(c)). GenMol adopts a masked discrete diffusion framework (Austin et al., 2021; Sahoo et al., 2024; Shi et al., 2024) with the BERT architecture (Devlin et al., 2019) to generate SAFE molecular sequences, thereby enjoying several advantages: (i) Discrete diffusion allows GenMol to exploit a molecular context that does not rely on the specific ordering of tokens and fragments by bidirectional attention. (ii) The non-autoregressive parallel decoding improves GenMol’s computational efficiency (Figure 2(b)). (iii) Discrete diffusion enables GenMol to explore chemical space with a simple yet effective remasking strategy. We propose fragment remasking (Figure 2(c6) and Figure 3), a strategy to optimize molecules by replacing certain fragments in a given molecule with masked tokens, from which diffusion generates new fragments. Utilizing fragments as the explorative unit instead of individual tokens is more in line with chemists’ intuition during optimization in drug discovery, and GenMol can effectively and efficiently explore the vast chemical space to find chemical optima. (iv) Discrete diffusion also makes it possible to apply guidance during generation based on the entire sequence. To this end, we propose molecular context guidance (MCG), a guidance method to improve the performance of GenMol by calibrating its predictions with information in a given molecular context.
@@
  • We introduce GenMol, a framework for unified and versatile molecule generation by building masked discrete diffusion that generates SAFE molecular sequences.
@@
  • We propose MCG, a guidance scheme for GenMol to effectively utilize molecular context information.
@@ 2 Related Work
There has been steady progress in applying discrete diffusion for discrete data generation, especially in NLP tasks (Hoogeboom et al., 2021; Austin et al., 2021; He et al., 2023; Zheng et al., 2024; Lou et al., 2024; Sahoo et al., 2024). This is mainly due to their non-autoregressive generation property, which leads to a potential for better modeling long-rang bidirectional dependencies and acclerating sampling speed, and their flexible design choices in training, sampling, and controllable generation (Sahoo et al., 2024). Notably, D3PM (Austin et al., 2021) introduced a general framework with a Markov forward process represented by transition matrices, and a transition matrix with an absorbing state corresponds to the masked language modeling (MLM) such as BERT (Devlin et al., 2019). Campbell et al. (2022) proposed a continuous-time framework for discrete diffusion models based on the continuous-time Markov chain (CTMC) theory. SEDD (Lou et al., 2024) introduced a denoising score entropy loss that extends score matching to discrete diffusion models. Sahoo et al. (2024) and Shi et al. (2024) proposed simple masked discrete diffusion frameworks, with the training objective being a weighted average of MLM losses across different diffusion time steps.
@@
Fragment-based molecular generative models refer to a class of methods that reassemble existing molecular substructures (i.e., fragments) to generate new molecules. They have been consider as an effective drug discovery approach as (i) assembling fragments simplifies the generation process and improves chemical validity and (ii) the unit that determines biochemical effect of a molecule is a fragment rather than an individual atom (Li, 2020). A line of works (Jin et al., 2020; Maziarz et al., 2021; Kong et al., 2022; Geng et al., 2023) used graph-based VAEs to generate novel molecules conditioned on discovered substructures. Xie et al. (2020) proposed to progressively add or delete fragments of molecular graphs using Markov chain Monte Carlo (MCMC) sampling. Yang et al. (2021) and Powers et al. (2023) used a reinforcement learning (RL) framework and classification, respectively, to progressively add fragments to the incomplete molecule. Graph-based genetic algorithms (GAs) (Jensen, 2019; Tripp & Hernández-Lobato, 2023) is a strong approach that decomposes parent molecules into fragments that are combined to generate an offspring molecule. However, since their generation is from random combinations of existing fragments with a local mutation of a small probability, they suffer from limited exploration in the chemical space. More recently, $f$-RAG (Lee et al., 2024a) introduced a fragment-level retrieval framework that augments the pre-trained molecular language model SAFE-GPT (Noutahi et al., 2024), where retrieving fragments from dynamically updated fragment vocabulary largely improves the exploration-exploitation trade-off. However, $f$-RAG still needs to train an information fusion module before adapting to various goal-oriented generation tasks.
@@ 3.1 Masked Diffusion
Masked diffusion models (Sahoo et al., 2024; Shi et al., 2024) are a simple and effective class of discrete diffusion models, and we follow MDLM (Sahoo et al., 2024) to define our masked diffusion. Formally, we define $\boldsymbol{x}$ as a sequence of $L$ tokens, each of which, denoted as $\boldsymbol{x}^{l}$, is a one-hot vector with $K$ categories (i.e., $\boldsymbol{x}^{l}_{i}\in\{0,1\}^{K}$ and $\sum_{i=1}^{K}\boldsymbol{x}^{l}_{i}=1$). Without loss of generality, we assume the $K$-th category represents the masking token, whose one-hot vector is denoted by m (i.e., $\textbf{m}_{K}=1$). We also define $\text{Cat}(\cdot;\boldsymbol{\pi})$ as a categorical distribution with a probability $\boldsymbol{\pi}\in\Delta^{K}$, where $\Delta^{K}$ represents the simplex over $K$ categories.
The forward masking process independently interpolates the probability mass between each token in clean data sequence $\boldsymbol{x}^{l}$ and the masking token m, defined as
$q(\boldsymbol{z}^{l}_{t}|\boldsymbol{x}^{l})=\text{Cat}(\boldsymbol{z}^{l}_{t};\alpha_{t}\boldsymbol{x}^{l}+(1-\alpha_{t})\textbf{m}),$ (1)
@@
$p_{\theta}(\boldsymbol{z}^{l}_{s}|\boldsymbol{z}^{l}_{t})=\begin{cases}\text{Cat}(\boldsymbol{z}^{l}_{s};\boldsymbol{z}^{l}_{t})&\boldsymbol{z}^{l}_{t}\neq\textbf{m}\\ \text{Cat}(\boldsymbol{z}^{l}_{s};\frac{(1-\alpha_{s})\textbf{m}+(\alpha_{s}-\alpha_{t})\boldsymbol{x}^{l}_{\theta}(\boldsymbol{z}_{t},t)}{1-\alpha_{t}})&\boldsymbol{z}^{l}_{t}=\textbf{m},\end{cases}$ (2)
where $\boldsymbol{x}_{\theta}(\boldsymbol{z}_{t},t)$ is a denoising network that takes the noisy data sequence $\boldsymbol{z}_{t}$ as input and predicts $L$ probability vectors for the clean data sequence. This parameterization designs the reverse process such that it does not change unmasked tokens. To train the denoising network $\boldsymbol{x}_{\theta}(\boldsymbol{z}_{t},t)$, the training objective, which implicitly approximates the negative ELBO (Sohl-Dickstein et al., 2015), is given by
@@ 3.2 SAFE Molecular Representation
Sequential Attachment-based Fragment Embedding (SAFE) (Noutahi et al., 2024) has been proposed to alleviate this problem. SAFE represents molecules as an unordered sequence of fragment blocks, thereby casting molecular design into sequence completion. SAFE is a non-canonical SMILES in which the arrangement of SMILES tokens corresponding to the same molecular fragment is consecutive. Molecules are decomposed into fragments by the BRICS algorithm (Degen et al., 2008) and the fragments are concatenated using a dot token (“.”) while preserving their attachment points. SAFE is permutation-invariant on fragments, i.e., the order of fragments within a SAFE string does not change the molecular identity.
@@ 4 Method
Figure 3: (a) Goal-directed hit generation and lead optimization process with GenMol. An initial fragment vocabulary is constructed by decomposing an existing molecular dataset (hit generation) or a seed molecule (lead optimization). Two fragments are randomly sampled from the vocabulary and attached, and GenMol performs fragment remasking. The fragment vocabulary is updated with the generated molecules for the next iteration. (b) Illustration of the molecular optimization trajectory with fragment remasking. With fragment remasking, GenMol can explore beyond the initial fragment vocabulary to find chemical optima.
We introduce GenMol, a universal molecule generation framework that can solve various drug discovery tasks. We first introduce the construction of the discrete diffusion framework on the SAFE representation in Section 4.1. Next, we describe the goal-oriented exploration strategy of GenMol with fragment remasking in Section 4.2. Finally, we describe MCG, a guidance scheme of GenMol by partially masking the given molecular context, in Section 4.3.
@@ 4.1 Masked Diffusion for Molecule Generation
We adopt masked discrete diffusion to generate SAFE sequences and establish a flexible and efficient molecule generation framework. Concretely, GenMol uses the BERT architecture (Devlin et al., 2019) as the denoising network $\boldsymbol{x}_{\theta}$ and the training scheme of MDLM. Utilizing discrete diffusion instead of an autoregressive model is more in line with the SAFE representation and has several advantages. First, due to the bidirectional attention in BERT, GenMol can leverage parallel decoding where all tokens are decoded simultaneously under discrete diffusion (Figure 5). As SAFE is fragment order-insensitive, this allows GenMol to predict masked tokens without relying on a specific ordering of generation while considering the entire molecule. The non-autoregressive parallel decoding scheme also improves GenMol’s efficiency. Furthermore, the discrete diffusion framework enables GenMol to explore the neighborhood of a given molecule with a remasking strategy.
At each masked index $l$, GenMol samples $\boldsymbol{z}^{l}_{s}$ based on the reverse unmasking process $p^{l}_{\theta}:=p_{\theta}(\boldsymbol{z}^{l}_{s}|\boldsymbol{z}_{t})$ specified by:
$p^{l}_{\theta,i}=\frac{\exp\left(\log\boldsymbol{x}^{l}_{\theta,i}(\boldsymbol{z}_{t},t)/\tau\right)}{\sum^{K}_{j=1}\exp\left(\log\boldsymbol{x}^{l}_{\theta,j}(\boldsymbol{z}_{t},t)/\tau\right)}\text{ for }i=1,\cdots,K,$ (4)
where $\log\boldsymbol{x}^{l}_{\theta,i}(\boldsymbol{z}_{t},t)$ is the logit predicted by the model and $\tau$ is the softmax temperature. All masked tokens are predicted in a parallel manner and GenMol confirms the top-$N$ confident predictions with additional randomness $r$ following Chang et al. (2022), where $N$ is the number of tokens to unmask at each time step. Trade-offs between molecular quality and diversity often arise in drug discovery, and GenMol can balance them through the softmax temperature $\tau$ and the randomness $r$. Further details about confidence-based sampling is provided in Section B.
4.2 Exploration in Chemical Space with GenMol
To perform goal-directed molecular optimization tasks, we propose a simple yet effective generation method (Figure 3) that consists of three steps: (1) fragment scoring, (2) fragment attaching, and (3) fragment remasking.
Fragment scoring.
We start with constructing a fragment vocabulary. A set of $D$ molecules $\mathcal{D}=\{\big{(}\boldsymbol{x}_{d},y(\boldsymbol{x}_{d})\big{)}\}^{D}_{d=1}$, where $y(\boldsymbol{x}_{d})$ is the target property of molecule $\boldsymbol{x}_{j}$, is decomposed into a set of $F$ fragments =$\mathcal{F}=\{\boldsymbol{f}_{k}\}_{k=1}^{F}$ using a predefined decomposition rule. We define the score of fragment $\boldsymbol{f}_{k}$ following Lee et al. (2024a) as:
$y(\boldsymbol{f}_{k})=\frac{1}{|\mathcal{S}(\boldsymbol{f}_{k})|}\sum_{\boldsymbol{x}\in\mathcal{S}(\boldsymbol{f}_{k})}y(\boldsymbol{x}),$ (5)
where $\mathcal{S}(\boldsymbol{f}_{k})=\{\boldsymbol{x}:\boldsymbol{f}_{k}\text{ is asubgraph of }\boldsymbol{x}\}$ and the top-$V$ fragments based on Eq. (5) are selected as the vocabulary.
Fragment attaching.
During generation, a molecule $\boldsymbol{x}_{\text{init}}$ is first generated by randomly selecting two fragments from the vocabulary and attaching them. Fragments or functional groups influence the chemical properties of a molecule and therefore, fragments that commonly occur in molecules with desirable properties are likely to carry them to new molecules. However, with fragment attaching alone, the model cannot generate new fragments that are not included in the initial vocabulary, resulting in suboptimal exploration in chemical space.
Fragment remasking.